### v1

@@ Abstract
The ACM Lifelog Search Challenge (LSC) is a venue that welcomes and compares systems that support the exploration of lifelog data, and in particular the retrieval of specific information, through an interactive competition format. This paper reviews the recent advances in interactive lifelog retrieval as demonstrated at the ACM LSC from 2022 to 2024. Through a detailed comparative analysis, we highlight key improvements across three main retrieval tasks: known-item search, question answering, and ad-hoc search. Our analysis identifies trends such as the widespread adoption of embedding-based retrieval methods (e.g., CLIP, BLIP), increased integration of large language models (LLMs) for conversational retrieval, and continued innovation in multimodal and collaborative search interfaces. We further discuss how specific retrieval techniques and user interface (UI) designs have impacted system performance, emphasizing the importance of balancing retrieval complexity with usability. Our findings indicate that embedding-driven approaches combined with LLMs show promise for lifelog retrieval systems. Likewise, improving UI design can enhance usability and efficiency. Additionally, we recommend reconsidering multi-instance system evaluations within the expert track to better manage variability in user familiarity and configuration effectiveness.
@@ I Introduction
Lifelogging [26] refers to the practice of recording aspects of daily life through wearable and other mobile devices, creating a rich and continuous record of personal experiences. As this wealth of data grows, so does the need for powerful retrieval systems that help users to search, explore, and make sense of their past. Whether for memory enhancement [10, 29], healthcare [69] or lifestyle monitoring [54, 92, 75], lifelog retrieval is key for subsequently transforming raw data into meaningful insights.
@@
Lifelog data presents unique retrieval challenges that differ from traditional multimedia search tasks such as those of TRECVID [8] or ImageCLEF [39], where search is asynchronous, and system performance can be evaluated in an automatic, streamlined manner. The first challenge is the sheer volume of data: Lifeloggers generate thousands of images, videos, audio recordings, locations, biometric signals and other sensor readings daily, making it difficult to locate specific events [26]. The second challenge is personal context: unlike traditional multimedia search, lifelog data is deeply personal, with unique emotions and memories attached to each record [36, 11, 16]. Retrieval systems must bridge this gap to provide meaningful results when used with lifelog data. The third challenge is multi-modality [17]; as mentioned above, lifelogs combine many different types of information, requiring retrieval systems to integrate diverse sources to be effective and comprehensive.
@@
Beyond these technical challenges, data availability remains a significant obstacle. Large-scale studies require extensive long-term datasets, yet finding lifeloggers willing to continuously record their daily lives, is difficult [22]. Consequently, only a handful of shared lifelog datasets exist [13], which means that researchers and system developers must search for events in data that are not their own. This lack of personal context makes it difficult to interpret events accurately. Without lived experience, searches may be entirely off, lacking the internal cues that lifeloggers naturally rely on. For example, a photo of a meal might appear in the dataset, but without additional context, was it a significant occasion or just an ordinary lunch?
@@
Another fundamental challenge is the nature of human memory, which is often fragmented, distorted, or shaped by emotions rather than factual details [12]. This makes query formulation particularly difficult. Coupled with the inherently noisy and unstructured nature of passively-collected lifelog data, this necessitates iterative, interactive search. This manner of search also closely mirrors how people naturally search for information—through trial and adjustment. Effective user interface (UI) and user experience (UX) design becomes crucial [15], allowing users to refine searches and explore their lifelogs intuitively and effectively, even without precise recall or first-hand experience of the events.
@@
Recognizing these challenges, the ACM Lifelog Search Challenge (LSC) was established to advance interactive lifelog retrieval systems, providing a competitive framework to evaluate their effectiveness in real-time scenarios. The ultimate aim is to move lifelog retrieval beyond basic data processing by creating tools that support reflection, true recall, and deeper engagement with personal histories. The LSC has been held annually since 2018, with the latest edition, LSC’24 [27], taking place at ACM ICMR 2024, in Phuket, Thailand. This paper reviews the key technical advances made in the LSC challenge over the last three years [28, 23, 27]. In particular, we provide an overview of the key developments observed across systems submitted to the past three LSC editions (2022–2024). We analyse trends in retrieval and interaction techniques, evaluate comparative system performance, and reflect on emerging research directions in lifelog search. Rather than introducing novel methods, this paper serves as a structured analysis of the field as seen through the lens of a recurring, shared evaluation challenge.
@@
The remainder of the paper is structured as follows: Section II describes the LSC competition setup, including the dataset, tasks, and evaluation procedures. Section III summarises the retrieval and interaction techniques employed by systems submitted to the past three editions of the challenge. In Section IV, we present a comparative analysis of system performance across tasks and years. Finally, Section V concludes with reflections on current limitations and outlines promising directions for future lifelog retrieval research.
@@ II The ACM Lifelog Search Challenge Workshops
LSC [28, 23, 27] is an annual challenge that is part of the ACM International Conference on Multimedia Retrieval (ICMR). The challenge was first introduced in 2018 and has attracted more participants than any other lifelogging challenge. The focus of the LSC is to evaluate the performance of interactive lifelog search engines in real time. Different systems compete with each other in a live environment, on-site or remotely; participants are given a set of queries and limited time to find the relevant images.
Over the years, both the types of tasks and the size of the datasets used have evolved, making the competition increasingly challenging. LSC’18 started with a single type of task, known-item search (KIS), where participants are given a textual query describing the content they need to find. The textual query, however, is not static; the organiser will provide an additional hint every 30 seconds (at 30, 60, 90, 120, and 150 seconds) to imitate a real-life scenario when someone gradually remembers more details over time. An example of LSC’s KIS task can be seen in Table I, and some images from the corresponding event in the dataset, which we refer to as ground truth, are shown in Figure II-A.
@@
For each KIS task, the participants are given a time limit of five minutes to find the relevant images. To solve the task, each participating team has to find a single relevant image for each KIS task and submit it to a host server [66]. The host server keeps track of time with a countdown clock and then assesses the submissions against the ground truth to evaluate their accuracy. In later years, further types of tasks were introduced, namely Ad-hoc and Question Answering (QA) tasks, which are described in Section II-B.
@@ II-A Datasets
Compared to our last analysis of LSC progress in 2021 [80], the LSC has since released a new and significantly expanded dataset for the 2022 and following campaigns. The dataset consists of 18 months of lifelog data (versus 85 days in the previous dataset [25]) from January 2019 to June 2020, making it one of the longest, richest and most comprehensive multimodal lifelog datasets available to the research community. It integrates images, biometrics, spatial and semantic data, enabling deep exploration of lifelog retrieval techniques. Key features of the dataset include:
  • Core image dataset: 725,000 first-person point-of-view lifelog images captured by a Narrative Clip device, fully redacted in 1024 $\times$ 768 resolution. Anonymisation has been applied by blurring faces and most readable text in a manual or semi-automated process. Private content that is not appropriate for public release is also removed. The images are stored in a hierarchical directory structure, with each directory representing a day and each image file named by its timestamp.
@@
  • Visual concepts: Two types of visual concepts are included in the metadata file: objects and scenes. The objects are detected using an object detection model trained on the Common Objects in Context (COCO) dataset widely used in computer vision research which has 80 object classes. [^1] For a full list of objects, please refer to https://github.com/amikelive/coco-labels/blob/master/coco-labels-2014_2017.txt. The scenes are detected using a scene recognition model trained on the Places365 dataset [94]. A total of 102 scene labels are detected, such as ‘waiting in line’, ‘working’, and ‘open area’. [^2] For a full list of scenes, please refer to https://github.com/CSAILVision/places365/blob/master/labels_sunattribute.txt The bounding boxes of the detected objects and confidence scores are provided for each detected concept. In addition, optical character recognition (OCR) outputs are also provided for the associated images.
@@
This dataset provides a rich source of information for lifelog retrieval research, enabling the development of advanced multimodal retrieval techniques. It is available upon request from the LSC organisers on the LSC website  [^3] http://lifelogsearch.org/lsc/2022/lsc_data/. All identifiable faces and personal documents in the dataset were blurred, masked, or omitted using a combination of automated and manual methods. The process of creating the dataset has been reviewed by Dublin City University ethics board and complies with GDPR principles. Ongoing efforts aim to balance research utility with the protection of participant privacy.
@@ II-B Tasks
As previously mentioned, the LSC has three types of tasks with textual queries: known-item search (KIS), ad-hoc search, and question answering (QA), see Table II. These tasks were created to reflect authentic information needs, occasionally including deliberate ambiguities to better simulate realistic search scenarios. For tasks other than KIS, complete ground truth is not created in advance, but human judges assess submissions as they arrive. In LSC’24, an assessment process was introduced to enhance the clarity of queries and consistency of relevance judgments. This process was adapted from the Video Browser Showdown (VBS) competition [9], and included a structured briefing session for reviewing and revising queries, along with a dry-run phase for testing open-set queries. These measures aimed to reduce ambiguity and support consistent, real-time judging by multiple assessors.
@@ II-B1 Known-Item Search
Known-Item Seach (KIS) is the oldest task in the LSC, where participants are given a textual query describing one or more scenes and have to find one of a set of relevant images within a time limit. The query is formulated as a series of temporally advancing descriptors revealed at 30-second intervals (see, e.g., Table I). The task mimics real-life scenarios in which people usually remember details slowly over time. There is no limit on the number of incorrect submissions a participant can make before a correct submission, which determines the resolution of the task for that participant. The time limit for this task is 300 seconds (5 minutes).
@@ II-B2 Ad-hoc Search
Ad-hoc search has been a common task in many retrieval benchmarking campaigns [49], including VBS [88], and the Japanese National Institute of Informatic NTCIR Lifelog Task [95]. However, it was only in 2022 that ad-hoc search tasks were introduced to the LSC. In these tasks, participants must submit as many images as possible relevant to the given textual query within a specified time limit. For example, a query used in the latest LSC competition was: ‘Find examples of when I was wearing my red Vietnam football shirt’. Because no ground truth is provided in advance, human judges are required to assess the relevance of the submissions. There is no predefined limit on the number of submissions (correct or incorrect) a participant can make, and the time limit is 180 seconds.
@@ II-B3 Question Answering
The idea of lifelog-based question answering was first explored through the LLQA benchmark [78], which demonstrated the potential of lifelog QA for supporting memory recall and highlighted its challenges in requiring reasoning across visual and temporal data. Around the same time, the QA task was introduced in LSC’22, initially requiring participants to retrieve an image that answered a given question. Since 2023, the task has evolved, now requiring participants to submit a text answer instead. For instance, an actual task from the LSC’24 competition was: ‘How many times did I visit an outdoor/farmers market in February 2020?’. Similar to ad-hoc search, the correctness of the answers is evaluated by human judges in real-time. Additionally, as in KIS tasks, there is no limit on the number of incorrect submissions a participant can make before a correct answer is submitted. Participants are given a time limit of 180 seconds to complete the task.
@@ II-C Synchronous Evaluation
If the task is not solved, the score is 0. Additionally, to discourage excessive submissions and to account for low precision in the submitted answers, the scoring formula includes a penalty of 10 points for each incorrect submission. This scoring formula combines correctness with a moderate time penalty, where submissions made at the five-minute limit receive 50% of the maximum score. This balance rewards both precision and efficiency while still crediting successful retrieval within the allowed time. To provide transparency, we also report correctness and time-to-answer metrics independently.
@@
where correct is the number of relevant images submitted, incorrect is the number of irrelevant images submitted, and total is the total number of relevant images aggregated over all submitting teams. The number of incorrect images is divided by 2 to reduce the penalty for wrong submissions because the number of submissions in the task is unlimited and the query hints are sometimes fuzzy. Similarly, if no relevant images are submitted, the score is 0.
@@ II-D Expert and Novice Users
In this paper, we focus on expert users because the performance of novice users exhibits high variability, making it difficult to draw consistent or meaningful conclusions. Novices may differ widely in their search strategies, prior knowledge, and ability to adapt to the system, leading to results that are harder to compare across participants. The limited time available for onboarding further affected their ability to engage with the systems consistently. While this limits conclusions about first-time usability, including novice results would have introduced noise and hindered fair comparison of system performance. By concentrating on expert users, we ensure a more controlled evaluation of retrieval performance, minimizing the impact of user-related variability.
However, this does not diminish the importance of UI and UX design in lifelog retrieval systems. Usability remains critical in real-world applications, as effective interfaces help users navigate complex lifelog data. While this study focuses on retrieval performance rather than usability testing, future work should incorporate novice users to assess the accessibility and user-friendliness of lifelog search systems.

### v2

@@ Abstract
The ACM Lifelog Search Challenge (LSC) is the only long-running benchmark that evaluates lifelog retrieval systems through real-time, human-in-the-loop interaction. This paper presents a longitudinal analysis of technical and interaction design progress at LSC from 2022 to 2024, conducted under fixed datasets, tasks, and evaluation protocols. Unlike prior surveys of lifelogging research, our study leverages competitive, synchronous evaluations to examine how retrieval architectures, interface design choices, and interaction strategies directly influence task performance. Through a comparative analysis of known-item search, question answering, and ad-hoc retrieval tasks, we identify a clear paradigm shift from concept-based pipelines toward embedding-driven and LLM-supported retrieval systems. We show how these shifts affect retrieval effectiveness, response time, and user behaviour, as well as provide empirical insight into how retrieval and interface design choices manifest in performance during interactive search. Beyond documenting emerging techniques, we expose evaluation artefacts such as performance variability across system instances and user familiarity effects, raising important considerations for future interactive benchmarks.
@@ I Introduction
Lifelogging [26] refers to the practice of recording aspects of daily life through wearable and other mobile devices, creating a rich and continuous record of personal experiences. As this wealth of data grows, so does the need for retrieval systems that help users to search, explore, and make sense of their past. Whether for memory enhancement [10, 29], healthcare [70] or lifestyle monitoring [54, 93, 76], lifelog retrieval is essential for subsequently transforming raw data into meaningful insights.
@@
Lifelog data presents unique retrieval challenges that distinguish it from traditional multimedia search tasks such as TRECVID [8] or ImageCLEF [39]. In those settings, search is asynchronous, and performance can be evaluated automatically. The first challenge is scale. Lifeloggers generate a high volume of images, video segments, and sensor readings daily, making it difficult to locate specific events [26]. The second challenge is personal context. Unlike conventional multimedia search, lifelog data is tied to personal memories and emotions [36, 11, 16]. Retrieval systems must bridge this gap to return meaningful results. The third challenge is multi-modality [17]. As mentioned above, lifelogs combine many different types of information, requiring retrieval systems to integrate diverse sources to be effective and comprehensive.
@@
Beyond these technical challenges, data availability remains a major obstacle. Large-scale studies require extensive long-term datasets, yet finding lifeloggers willing to record their daily lives presents major recruitment and retention challenges [22]. Consequently, only a handful of shared lifelog datasets exist [13], forcing researchers to search for events in data that are not their own. This lack of personal context makes it difficult to accurately interpret events. Without lived experience, searches may be substantially inaccurate or contextually irrelevant, lacking the internal cues that lifeloggers naturally rely on. For example, a photo of a meal might appear in the dataset, but without additional context, was it a significant occasion or just an ordinary lunch?
@@
Another fundamental challenge is the nature of human memory, which is often fragmented, distorted, or shaped by emotions rather than factual details [12]. This makes query formulation particularly difficult. Combined with the noise and unstructured nature of passively collected data, this makes iterative, interactive search essential. This manner of search also closely mirrors how people naturally search for information – through trial and adjustment. Effective user interface (UI) and user experience (UX) design becomes crucial [15]. Good interfaces allow users to refine searches and explore their lifelogs intuitively, even without precise recall or first-hand experience.
@@
To address these challenges, the ACM Lifelog Search Challenge (LSC) was established to advance interactive lifelog retrieval through a competitive, real-time evaluation framework The ultimate goal is to move lifelog retrieval beyond basic data processing by creating tools that support reflection, true recall, and deeper engagement with personal histories. The LSC has been held annually since 2018, with the latest edition, LSC’24 [27], taking place at ACM ICMR 2024, in Phuket, Thailand. This paper reviews the key technical advances demonstrated in the past three LSC editions (2022–2024) [28, 23, 27]. We analyse trends in retrieval and interaction techniques, evaluate comparative system performance, and reflect on emerging research directions in lifelog search. Rather than proposing novel methods, this paper provides a structured analysis of the field through the lens of a recurring, shared evaluation challenge.
Although this study focuses on the LSC series, its longitudinal design across identical datasets and tasks provides a uniquely controlled lens for observing algorithmic and interaction-design evolution. Comparable long-term, human-in-the-loop benchmarks are rare in multimedia retrieval. Analysing LSC systems, therefore, offers insight into how the broader field transitions from concept-based to embedding- and LLM-driven paradigms under identical experimental conditions.
@@
The remainder of the paper is structured as follows: Section II describes the LSC competition setup, including the dataset, tasks, and evaluation procedures. Section III summarises the retrieval and interaction techniques employed by the systems submitted to the past three editions of the challenge. In Section IV, we present a comparative analysis of system performance across tasks and years. Finally, Section V concludes with reflections on current limitations and outlines promising directions for future lifelog retrieval research.
@@ II The ACM Lifelog Search Challenge Workshops
LSC [28, 23, 27] is an annual challenge that is part of the ACM International Conference on Multimedia Retrieval (ICMR). The challenge was first introduced in 2018 and has attracted the highest participation among similar lifelogging challenges. The focus of the LSC is to evaluate the performance of interactive lifelog search engines synchronously. Different systems compete with each other in a live environment (on-site or remote); participants are given a set of queries and a limited time to find the relevant images.
Over the years, both the types of tasks and the size of the datasets used have evolved, making the competition increasingly challenging.
LSC’18 started with a single type of task, known-item search (KIS), where participants are given a textual query describing the content they need to find. This query advances in stages to imitate gradual memory recall (see Section II-B1 and Table I). Participants submit a single relevant image to the host server [67], which manages the time limit and assesses accuracy against the ground truth. Some images from the corresponding event in the dataset, which we refer to as ground truth, are shown in Figure II-A. In subsequent years, additional task types were introduced, specifically Ad-hoc and Question Answering (QA) tasks, which are described in Section II-B.
@@ II-A Datasets
Building upon our previous analysis of LSC progress in 2021 [81], the LSC introduced a new and substantially larger dataset for the 2022 and subsequent campaigns. The dataset consists of 18 months of lifelog data (a substantial increase over the 85 days in the preceding dataset [25]) from January 2019 to June 2020, making it one of the longest, richest and most comprehensive multimodal lifelog datasets available to the research community. It integrates images, biometrics, spatial and semantic data, enabling deep exploration of lifelog retrieval techniques. Key features of the dataset include:
  • Core image dataset: 725,000 first-person point-of-view lifelog images captured by a Narrative Clip device, fully redacted in 1024 $\times$ 768 resolution. Anonymisation was applied through a combination of manual and semi-automated blurring of faces and legible text. Private content that is not appropriate for public release is also removed. The images are stored in a hierarchical directory structure, with each directory representing a day and each image file named by its timestamp.
@@
  • Visual concepts: Two types of visual concepts are included in the metadata file: objects and scenes. The objects are detected using an object detection model trained on the Common Objects in Context (COCO) dataset, a standard resource in computer vision research, which has 80 object classes. [^1] For a full list of objects, please refer to https://github.com/amikelive/coco-labels/blob/master/coco-labels-2014_2017.txt. The scenes are detected using a scene recognition model trained on the Places365 dataset [95]. A total of 102 scene labels are detected, such as ‘waiting in line’, ‘working’, and ‘open area’. [^2] For a full list of scenes, please refer to https://github.com/CSAILVision/places365/blob/master/labels_sunattribute.txt The bounding boxes of the detected objects and confidence scores are provided for each detected concept. In addition, optical character recognition (OCR) outputs are also provided for the associated images.
@@
This dataset provides a valuable resource of information for lifelog retrieval research, enabling the development of advanced multimodal retrieval techniques. It is available upon request from the LSC organisers on the LSC website  [^3] http://lifelogsearch.org/lsc/2022/lsc_data/. All identifiable faces and personal documents in the dataset were blurred, masked, or omitted using a combination of automated and manual methods. The process of creating the dataset has been reviewed by the Dublin City University ethics board and complies with GDPR principles. Ongoing efforts aim to balance research utility with the protection of participant privacy.
@@ II-B Tasks
As previously mentioned, the LSC has three types of tasks with textual queries: known-item search (KIS), ad-hoc search, and question answering (QA), see Table II. These tasks were created to reflect authentic information needs, often incorporating deliberate ambiguities to closely simulate realistic search scenarios. For tasks other than KIS, a complete ground truth is not created in advance, but human judges assess submissions in real-time. In LSC’24, an assessment process was introduced to enhance the clarity of queries and consistency of relevance judgments. This process was adapted from the Video Browser Showdown (VBS) competition [9], and included a structured briefing session for query review and revision, alongside a dry-run phase for testing open-set queries. These measures aimed to minimise ambiguity and ensure consistent, synchronous relevance judgments by multiple assessors.
@@ II-B1 Known-Item Search
Known-Item Search (KIS) is the oldest task in the LSC, where participants are given a textual query describing one or more scenes and are required to retrieve one of a predefined set of relevant images within a specified time limit. The query is formulated as a series of temporally advancing descriptors revealed at 30-second intervals (see, e.g., Table I). The task simulates real-life scenarios where individuals gradually recall details over time. There is no limit on the number of incorrect submissions a participant can make before a correct submission, which determines the successful completion of the task for that participant. The time limit for this task is 300 seconds (5 minutes).
@@ II-B2 Ad-hoc Search
Ad-hoc search has been a common task in many retrieval benchmarking campaigns [49], including VBS [89], and the Japanese National Institute of Informatics NTCIR Lifelog Task [96]. However, it was only in 2022 that ad-hoc search tasks were introduced to the LSC. In these tasks, participants must submit as many images as possible relevant to the given textual query within a specified time limit. For example, a query used in the latest LSC competition was: ‘Find examples of when I was wearing my red Vietnam football shirt’. Because no ground truth is provided in advance, human judges are required to assess the relevance of the submissions. There is no predefined limit on the number of submissions (correct or incorrect) a participant can make, and the time limit is 180 seconds.
@@ II-B3 Question Answering
The concept of lifelog-based question answering was initially explored through the LLQA benchmark [79], which demonstrated the potential of lifelog QA for supporting memory recall and underscored its challenges in requiring reasoning across visual and temporal data. Concurrently, the QA task was introduced in LSC’22, initially requiring participants to retrieve an image that answered a given question. Since 2023, the task has evolved, now requiring participants to submit a text answer instead. For instance, an actual task from the LSC’24 competition was: ‘How many times did I visit an outdoor/farmers market in February 2020?’. Similar to ad-hoc search, the correctness of the answers is evaluated by human judges in real-time. Additionally, as in KIS tasks, there is no limit on the number of incorrect submissions a participant can make before a correct answer is submitted. Participants are given a time limit of 180 seconds to complete the task.
@@ II-C Synchronous Evaluation
If the task is not solved, the score is 0. Additionally, to discourage excessive submissions and to account for low precision in the submitted answers, the scoring formula includes a penalty of 10 points for each incorrect submission. This balance rewards both precision and efficiency, with successful submissions made at the five-minute time limit receiving 50% of the maximum score. To provide transparency, we also report correctness and time-to-answer metrics independently.
@@
where correct is the number of relevant images submitted, incorrect is the number of irrelevant images submitted, and total is the total number of relevant images aggregated over all submitting teams. The number of incorrect images is divided by 2 to reduce the penalty for wrong submissions because the number of submissions in the task is unlimited, and the query hints are sometimes fuzzy. Similarly, if no relevant images are submitted, the score is 0.
@@ II-D Expert and Novice Users
In this paper, we focus on expert users because the performance of novice users exhibits high variability, making it difficult to draw consistent or meaningful conclusions. This variability stems from differences in search strategies and system familiarity, which is compounded by limited onboarding time, making results harder to compare across participants. While this limits conclusions about first-time usability, concentrating on expert users ensures a more controlled evaluation of retrieval performance, minimising the impact of user-related variability and noise for fair system comparison.
However, the importance of UI and UX design is not diminished, as effective interfaces remain critical for users to navigate complex lifelog data in real-world applications. While this study focuses on retrieval performance rather than usability testing, future work should incorporate novice users to assess the accessibility and user-friendliness of lifelog search systems.
II-E Relation to Prior Benchmarks and Research
LSC shares lineage with prior interactive retrieval benchmarks such as TRECVID[8], VBS[89], ImageCLEF Lifelog[57], and NTCIR-Lifelog[97], yet differs in its focus and evaluation methodology. While TRECVID and VBS primarily benchmark large-scale video retrieval and interactive search efficiency in general-purpose datasets, lifelogging tasks at ImageCLEF and NTCIR introduced retrieval within personal or daily-life contexts. However, both ImageCLEF and NTCIR Lifelog tasks follow asynchronous evaluation protocols, where systems are assessed offline on pre-defined queries and ranked outputs. In contrast, the Lifelog Search Challenge (LSC) adopts a real-time, interactive format in which human users actively search through lifelog data using live systems. This unique approach complements earlier benchmarks by integrating algorithmic retrieval quality with user interaction, interface design, and time-constrained performance, enabling the study of human-in-the-loop search behaviour and system responsiveness under authentic, real-world conditions.